OpenStudy (anonymous):

Prove {2 * sin (theta) * cos(theta) - cos(theta)}/{1 - sin(theta) + sin^2(theta) - cos^2(theta)} = cot(theta)

OpenStudy (anonymous):

madhur.. u wanna work on this problem

OpenStudy (anonymous):

yaa

OpenStudy (anonymous):

do u know the formula for sin2theta have u tried tht?

OpenStudy (anonymous):

these things can be tricky, often a good idea to look for clues in the question. In this case, the right hand side is cot or 1/Tan = Cos/Sin so if you could somehow get the left hand side to be cos/sin you would be home.

OpenStudy (anonymous):

I've got a Idea ... U can take cos common out of the numerator and convert 1 into sin^2 + cos^2

OpenStudy (anonymous):

I'm talking about the 1 in denominator

OpenStudy (anonymous):

cos^2 gets cancel and then u take sin common and u get ur answer :)

OpenStudy (anonymous):

Him is tht right? I think It is

OpenStudy (anonymous):

Looks pretty good to me...

myininaya (myininaya):

that looks absolutely good so just looking at the bottom: 1=sin^2x+cos^2x then use that cos^2x-cos^2x=0 so you have on bottom sin^2x+sin^2x-sinx=2sin^2x-sinx=sinx(2sinx-1) and on top you have cosx(2sinx-1)